please help me i am not able to help my daughter with this 6th grade math question. can you tell me how much materical you would need to build a shadow box with these dimensions? there is no front to the box to display items.
8 inches high, 7 inches in length and 5 inches deep.

Answer by ankor@dixie-net.com(22740)

how much material you would need to build a shadow box with these dimensions? there is no front to the box to display items.
8 inches high, 7 inches in length and 5 inches deep.
:
Surface area of a box with 5 sides
S.A. = (8*7) + 2(8*5) + 2(7*5)
S.A. = 56 + 2(40) + 2(35)
S.A. = 56 + 80 + 70
S.A. = 206 sq/in